OpenStudy (anonymous):

solve (t^(2)+25)(dx/dt)=(x^(2)+36)

OpenStudy (anonymous):

Separate the variables to get \[\begin{align*} \frac{dx}{x^2+36}=\frac{dt}{t^2+25} \end{align*}\] For both integrals, you can make the appropriate trig sub, \((\text{ind var})=\sqrt{\text{constant}}~\tan(\text{new var})\).

OpenStudy (anonymous):

In other words, as an example, set \(x=6\tan a\), then \(dx=6\sec^2a~da\), and the left side becomes \[\int\frac{6\sec^2a}{36\tan^2a+36}~da\]

OpenStudy (anonymous):

okay but i need to solve for x

OpenStudy (anonymous):

Alright, so what's you're implicit solution?

OpenStudy (anonymous):

Left hand side: \[\begin{align*}\frac{1}{6}\int\frac{\sec^2a}{\tan^2a+1}~da&=\frac{1}{6}\int\frac{\sec^2a}{\sec^2a}~da\\\\ &=\frac{1}{6}\int da\\\\ &=\frac{1}{6}a\color{gray}{+C}\\\\ &=\frac{1}{6}\tan^{-1}\frac{x}{6}\color{gray}{+C} \end{align*}\] What's the right side going to be?

OpenStudy (anonymous):

(1/5)arctan(x/5)+C

OpenStudy (anonymous):

Well, x should be t, but yes, that's right.

OpenStudy (anonymous):

So you have \[\frac{1}{6}\tan^{-1}\frac{x}{6}=\frac{1}{5}\tan^{-1}\frac{t}{5}+C\] Multiply both sides by 6: \[\tan^{-1}\frac{x}{6}=\frac{6}{5}\tan^{-1}\frac{t}{5}+C\] Take the tangent of both sides: \[\frac{x}{6}=\tan\left(\frac{6}{5}\tan^{-1}\frac{t}{5}+C\right)\] You can use an identity for the right side: \[\tan(\alpha\pm\beta)=\frac{\tan\alpha\pm\tan\beta}{1\mp\tan\alpha\tan\beta}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Correct